OpenStudy (boldgeorge):

Need quick help with a graph! Medal+Fan! Minutes, x|Cost,y ($) 01000 | y=65 a) Draw the graph of this piecewise linear function. b) Calculate the cost of using the Internet for 400 mins, 850 min, and 18 H.

OpenStudy (anonymous):

Graph starts at x = 0 from y = 15, and goes on till x = 600 (from x = 0 to x = 600 it's a straight line y = 15, parallel to the x axis. From x = 600 - 1000, it's a straight line with gradient 1/2 and intercept 285 ( y = 15 + 0.5(x-600) = > y = 0.5 x +15 - 300 = .5x - 285. From x = 100-, it's again a straight line parallel to x axis, y = 65.

OpenStudy (boldgeorge):

@NNL1991 I'm having trouble actually making this into a line graph.. would you be able to help? Or anyone else?